OpenStudy (anonymous):

that website is complicating.... what is 1/3(5n+12)+1/3n=-1

OpenStudy (anonymous):

Again, multiply both sides by 3.

OpenStudy (anonymous):

what do i multiply 3 on the left side?

OpenStudy (anonymous):

\[1/3(5n+12)+(n/3)=-1\]

OpenStudy (anonymous):

The whole left side: \[3(\frac{1}{3}(5n + 12) + \frac{1}{3}n) = 3(-1)\]

OpenStudy (anonymous):

Then you will of course need to distribute the 3, but it just cancels with all the fractions, so it's not hard.

OpenStudy (anonymous):

\[(5n+12+n+3)/3 = -1\]

OpenStudy (anonymous):

shouldnt i try to solve it first? and why did math pro change 1/3 to n/3?

OpenStudy (anonymous):

Any time you perform an operation (like adding, multiplying) on an equation that isn't already part of the equation you have to do it to every term on both sides of the equation.

OpenStudy (anonymous):

He changed \(\frac{1}{3}n\) to \(\frac{n}{3}\), since that's how fractions are multiplied.

OpenStudy (anonymous):

now multiply 3 on both sides\[3[(5n+12+n+3)/3] = -1*3\]

OpenStudy (anonymous):

It's the same as \(\frac{(1)(n)}{3}\), but the 1 as a coefficient is not necessary, and is normally not shown in equations.

OpenStudy (anonymous):

3(6n+15)=-3?

OpenStudy (anonymous):

Very close.

OpenStudy (anonymous):

But the 3 out front cancelled with the 1/3, so it's just 6n + 16 = -3

OpenStudy (anonymous):

elfigo you forgot one thing

OpenStudy (anonymous):

err 6n+15=-3 rather.

OpenStudy (anonymous):

yea :D

OpenStudy (anonymous):

So now you just have to solve for n. You can do this by first subtracting 15 from both sides, then dividing both sides by the coefficient on the n variable (6)

OpenStudy (anonymous):

sorry my intwernet kicked me off where were we at

OpenStudy (anonymous):

@elfigo 6n+15=-3 now subtract 15 on the sides 6n+15-15=-3-15